OpenStudy (anonymous):

A baseball is thrown up into the air. The ball's height in feet is modeled by the equation h(t) = -16t2 + 32t + 10. (h = height, t = time in seconds) Find the maximum height of the ball. a) 10 feet b) 2.3 feet c) 26 feet d) 23 feet ?

OpenStudy (anonymous):

If you find the derivative of the function: h'(t)=-32t+32 This shows the rate of change of the function. So when it is =0 it means that it is a min/max of the function. (here it is a max) So: -32t+32=0 solving gives t=1 (at t=1 h(t) has max) Put it back in the function: -16+32+10=26 So the answer is C.